If the base of a ladder is placed on the ground 4ft from a house, how tall must the ladder be (to the nearest foot) to reach the chimney top that extends 18.5ft above the ground?
Answer by richwmiller(17219)

we are setting up a right triangle.
a^2+b^2=c^2
where b=4 ft
and a=18.5 ft.
4^2+18.5^2=c^2
where c is the length of the ladder
and of course we only want positive numbers.
Remember we want to the nearest foot.